OpenStudy (idealist10):

Solve the initial-value problem y'-2y=xy^3, y(0)=2sqrt(2).

OpenStudy (idealist10):

I did Bernoulli method and got y=+/-sqrt(4/x-16+(e^(4x))/C) and when I plug in 0 into x, I got undefined in 4/x.

OpenStudy (amistre64):

let z = y^-2, z' = -2y^-3 y' y^-3 y' - 2y^-2 = x -2y^-3 y' + 4y^-2 = -2x z' + 4z = -2x does this look correct?

OpenStudy (idealist10):

Yes, I got v'+4v=-2x, I did v instead of z like yours. And then the integrating factor is e^4x, right?

OpenStudy (amistre64):

yes

OpenStudy (idealist10):

Then I got (e^4x)v'+4ve^4x=-2xe^4x e^(4x)v=-2*integral of xe^(4x) dx I did integration by parts and got (xe^4x)/4-(e^4x)/16+C on the right side.

OpenStudy (amistre64):

well, solve the homogenous z' + 4z = 0 e^(4z)z' + 4e^(4z) z = 0 e^(4z) z = C z = Ce^(-4z) we solve it in z right?

OpenStudy (idealist10):

But why did you solve for z like that?

OpenStudy (amistre64):

y' + y = 0 e^x y implicit is x'e^x y + e^x y' just trying to recollect is all, assuming dy/dx, then x'=1 i spose thats correct :)

OpenStudy (amistre64):

z = Ce^(-4x) therefore y^-2 = Ce^(-4x) (1/y)^2 = Ce^(-4x) 1/y = Ce^(-2x) y = Ce^(2x) does that look valid?

OpenStudy (idealist10):

Let me check the answer.

OpenStudy (amistre64):

the wolf is showing me e^(2x) is in the answer

OpenStudy (idealist10):

The answer in the book says y=(2sqrt(2))/(sqrt(1-4x)).

OpenStudy (amistre64):

is there an i in there someplace?

OpenStudy (idealist10):

Well, the answer in the book doesn't have i, but if you plug in a number like 4 for example into x, the answer will give you a non-real result if you evaluate it in a graphing calculator.

OpenStudy (amistre64):

trying to recall a wronskian at the moment, are at least variation of parameters

OpenStudy (idealist10):

@myininaya

OpenStudy (idealist10):

@satellite73 @Hero @campbell_st

OpenStudy (idealist10):

@ganeshie8

OpenStudy (anonymous):

\[y'-2y=xy^3\] divide by y^3 \[y ^{-3}y \prime-2 y ^{-2}=x\] \[put~ y ^{-2}=t,-2 y ^{-3}y \prime=\frac{ dt }{ dx }\] \[y ^{-3}y \prime=-\frac{ 1 }{ 2 }\frac{ dt }{ dx }\] \[-\frac{ 1 }{ 2 }\frac{ dt }{ dx }-2 t=x\] \[\frac{ dt }{ dx }+4t=-2x\] \[I.F.=e ^{\int\limits 4 dx}=e ^{4x}\] c.s. is \[t~e ^{4x}=\int\limits -2x~e ^{4x}dx+c\] \[y ^{-2}e ^{4x}=-2~x \frac{ e ^{4x} }{ 4 }-\int\limits (-2)\frac{ e ^{4x} }{ 4 }dx+c\] \[=-\frac{ e ^{4x} }{ 2 }+\frac{ e ^{4x} }{ 2*4 }+c\]

OpenStudy (anonymous):

when x=0,y=2sqrt2, find c and then substitute.